How is the zero product property used to solve a quadratic equations? What would be an good example of this?
Answer by richard1234(7193)

Sometimes if a quadratic, for example x^2 - 5x + 4 = 0 factors to (x-4)(x-1) = 0, we can deduce that the roots occur when x-4 = 0 or x-1 = 0, i.e. x = 4 or 1.